# PoisonedFL: Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v1

@@ Abstract.
doi: XXXXXXX.XXXXXXX
1. Introduction
Federated Learning (FL) enables clients–such as smartphones, self-driving cars, and IoT devices–to collaboratively train a model without sharing their raw local training data with a central server (Konečnỳ et al., 2016; McMahan et al., 2017). Typically, FL involves multiple training rounds, each of which consists of three key steps: 1) the server sends the current global model to the clients or a subset of them; 2) a client trains a local model based on the current global model and its local training data, and sends the model update (i.e., the difference between local model and global model) to the server; and 3) the server aggregates the clients’ model updates according to an aggregation rule and adds the aggregated model update to the global model. FL has wide applications across diverse domains, including healthcare (mel, [n. d.]; Kairouz et al., 2021; Rieke et al., 2020) and finance (web, [n. d.]; Yang et al., 2019).
However, due to the distributed nature, FL is fundamentally vulnerable to model poisoning attacks (Baruch et al., 2019; Cao and Gong, 2022; Fang et al., 2020; Shejwalkar and Houmansadr, 2021). In particular, an attacker can control some malicious clients and send carefully crafted malicious model updates to the server to poison the global model. The malicious clients could be compromised genuine clients or injected fake clients. The poisoned global model misclassifies many inputs indiscriminately, i.e., it has a large testing error rate (also known as untargeted model poisoning attacks).
Existing model poisoning attacks suffer from two key limitations. First, they (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021; Cao and Gong, 2022) only leverage in-round consistency, i.e., they send the same or similar malicious model updates in a single training round to amplify the attack effect, but neglect multi-round consistency. For instance, Fang attack (Fang et al., 2020) crafts the malicious model updates in a given training round to maximally deviate the global model in that round, while MPAF (Cao and Gong, 2022) sets the malicious model updates in a training round as the magnified difference between the current global model and a random model. Such a strategy limits the attack effectiveness when defenses are deployed. This is because defenses weaken the attack effect in every training round by filtering and clipping the malicious model updates; and the weakened attack effects further self-cancel across training rounds due to the lack of multi-round consistency, as shown by our measurement results in Section 4. Second, most existing attacks (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021) require knowledge of genuine clients’ model updates or local training data. They often assume compromising a significant number of genuine clients to obtain such knowledge, which is notably impractical. These attacks with suboptimal effectiveness and/or impractical assumptions give a false sense of FL security.
@@ 1. Introduction
We evaluate our PoisonedFL on five benchmark datasets. Our results show that, under the least knowledge scenario – that is, no knowledge about the defense nor genuine local training data/models – PoisonedFL can break eight state-of-the-art FL defenses (Blanchard et al., 2017; Cao et al., 2021a, 2022; Nguyen et al., 2022; Sun et al., 2019; Yin et al., 2018; Zhang et al., 2022). For instance, PoisonedFL can reduce the global models to random guessing in many cases. Moreover, in most cases, PoisonedFL achieves higher testing error rates than existing attacks (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021; Cao and Gong, 2022) when a defense is deployed even if we assume the attacker has access to the local models of all genuine clients. In the remaining cases, both PoisonedFL and some existing attacks reduce the global models to random guessing. Furthermore, PoisonedFL remains effective across different FL settings including the fraction of malicious clients, the degree of non-IID of the genuine clients’ local training data, the number of epochs used to train a local model, and the fraction of clients selected to participate in each training round.
@@
2. Background and Related Work
2.1. Federated Learning (FL)
@@
Federated Learning (FL) (Konečnỳ et al., 2016; McMahan et al., 2017) is a machine learning paradigm that allows many clients (e.g., smartphones) to collaboratively train a model without sending their raw local training data to a server. Suppose we have $n$ clients and each client has a local training dataset $\mathcal{D}_{i}$, $i=1,2,\cdots,n$. FL aims to solve an optimization problem that is defined as $\min\sum_{i=1}^{n}\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$, where $\bm{w}\in\mathbb{R}^{d}$ is the model (called global model), $d$ is the number of parameters in the model $\bm{w}$, and $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ is the local loss function for client $i$. FL iteratively trains the global model $\bm{w}$. In the $t$-th training round, the server first distributes the current global model $\bm{w}^{t-1}$ to all clients or a subset of them. For simplicity, we assume all clients are selected, but we will also evaluate a subset of them are selected in each training round in our experiments. Each client $i$ then trains a local model $\bm{w}_{i}^{t}$ based on the current global model and its local training data. For instance, client $i$ initializes $\bm{w}_{i}^{t}$ as $\bm{w}^{t-1}$ and iteratively updates $\bm{w}_{i}^{t}$ via minimizing $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ using Stochastic Gradient Descent (SGD). After training the local model $\bm{w}_{i}^{t}$, client $i$ sends the model update $\bm{g}^{t}_{i}=\bm{w}^{t}_{i}-\bm{w}^{t-1}$ to the server. Finally, the server aggregates the received model updates and updates the global model as follows:
(1) $\bm{w}^{t}=\bm{w}^{t-1}+\bm{g}^{t},$
@@
where $\bm{g}^{t}={{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the aggregated model update, ${AR}$ denotes the aggregation rule used by the server, and $\{\bm{g}^{t}_{i}\}_{i\in[1,n]}$ is the set of clients’ model updates. For instance, when ${AR}$ is FedAvg (McMahan et al., 2017), the aggregated model update ${{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the average of the clients’ model updates.
2.2. Poisoning Attacks to FL
@@
In model poisoning attacks (Bagdasaryan et al., 2020; Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021), malicious clients corrupt the global model via sending carefully crafted malicious model updates to the server. In a targeted attack (Bagdasaryan et al., 2020), a poisoned global model only misclassifies attacker-chosen inputs with certain properties (e.g., embedded with an attacker-chosen trigger). In contrast, in an untargeted attack (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021), a poisoned global model indiscriminately misclassifies a large number of clean inputs, leading to a large testing error rate and eventually denial-of-service attacks. Therefore, we focus on untargeted attacks in this work.
@@
Requiring genuine clients’ information:  One category of model poisoning attacks (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021) require that an attacker has access to genuine clients’ model updates or local training data. To obtain such knowledge, these attacks often assume an attacker can compromise a large fraction of genuine clients. For instance, in a little is enough (LIE) attack (Baruch et al., 2019), the attacker computes the average of the genuine model updates of the compromised genuine clients and adds perturbation to them to craft the malicious model updates in each training round. Concurrent to LIE, Fang et al. (Fang et al., 2020) proposed the first optimization-based attack framework that can be tailored to any aggregation rule. Their key idea is that the attacker crafts the malicious model updates such that the aggregated model update after attack deviates substantially from the before-attack one in each training round. Building upon this attack framework (Fang et al., 2020), the authors in (Shejwalkar and Houmansadr, 2021) assume the malicious model update to be the sum of the before-attack aggregated model update and a fixed perturbation vector scaled by a factor, where the scaling factor is optimized to maximize the deviation between the after-attack aggregated model update and the before-attack one.
@@
Not requiring genuine clients’ information:  The other category of model poisoning attacks (Cao and Gong, 2022) do not rely on any genuine clients’ information. Therefore, they only assume that an attacker can inject fake clients into an FL system, which is more realistic than compromising genuine clients. Since genuine clients’ information is not available, these attacks craft malicious model updates based on heuristics. For instance, in MPAF (Cao and Gong, 2022), the attacker crafts malicious model updates such that the global model progressively moves towards a random model in each training round.
@@
Limitations of existing attacks:  Existing attacks suffer from the following limitations. First, the first category of attacks require an attacker to compromise a large fraction of genuine clients, which is not affordable in production FL (Shejwalkar et al., 2022). Second, both categories of attacks craft malicious model updates independently in different training rounds, leading to self-cancellation of attack effect across training rounds. As a result, they achieve suboptimal attack effectiveness when defenses are deployed.
2.3. Defenses against Poisoning Attacks
@@
Byzantine-robust aggregation rules:  Instead of using the average of the clients’ model updates to update the global model, Byzantine-robust aggregation rules (Blanchard et al., 2017; Cao et al., 2021a; Mhamdi et al., 2018; Nguyen et al., 2022; Sun et al., 2019; Yin et al., 2018; Fang et al., 2022; Xu et al., 2024) typically employ statistical analysis to filter out or clip outlier model updates before aggregating them. For instance, Multi-Krum (Blanchard et al., 2017) iteratively selects the model update that is the closest to its nearest neighbor model updates, and aggregates them. Dimension-wise Median (or Trimmed Mean) (Yin et al., 2018) uses the median (or trimmed-mean) of the clients’ local models for each dimension. Norm Bound (Sun et al., 2019) aims to minimize the impact of potential malicious model updates by clipping the magnitudes of all model updates. FLTrust (Cao et al., 2021a) assumes the server uses a clean and small dataset to bootstrap trust. Specifically, in FLTrust (Cao et al., 2021a), the server trains a server model and assigns trust scores to clients by measuring the similarity between the server model update and clients’ model updates. Under some assumptions about the learning problem and clients’ local training data distribution, these Byzantine-robust aggregation rules often can theoretically guarantee that the learnt global model does not change much when the number of malicious clients is bounded. However, they are still vulnerable to our attack because these assumptions often do not hold and a bounded change in the parameters of the global model does not guarantee a small change of error rate.
@@
Provably robust defenses:  Provably robust defenses (Cao et al., 2021b, 2022; Xie et al., 2021) can provide stronger theoretical guarantees than Byzantine-robust aggregation rules. In particular, given a number of malicious clients, a provably robust defense can guarantee a lower bound of testing accuracy no matter what poisoning attacks the malicious clients use. For instance, FLCert (Cao et al., 2022) trains multiple global models, each of which is trained using a subset of clients. Given a testing input, FLCert predicts its label by taking a majority vote among the multiple global models. A major limitation of these provably robust defenses is that they can only tolerate a small number of malicious clients, i.e., the lower bound of testing accuracy reduces to 0 when the number of malicious clients is moderate. As a result, they are still vulnerable to our attack since an attacker can easily inject a large number of fake clients.
@@
Malicious clients detection:  These defenses (Li et al., 2020; Zhang et al., 2022) aim to detect malicious clients during or after the training process. After detecting malicious clients, the server can further efficiently recover/re-train an accurate global model using the remaining clients (Cao et al., 2023). For instance, FLDetector (Zhang et al., 2022) aims to detect malicious clients by utilizing the inconsistency of a client’s model updates across multiple training rounds. Their key idea is that the model updates of a benign client in multiple training rounds are all calculated based on the client’s local training data, while the model updates of a malicious client are crafted without following the learning procedure. As a result, the model updates of a benign client satisfy certain consistency constraints across training rounds, while those of a malicious client do not, making it possible to distinguish between benign and malicious clients. In our attack, we enforce the malicious model updates of a malicious client to be consistent across training rounds, making FLDetector ineffective.
3. Threat Model
@@
Attacker’s goal:  Following previous works  (Baruch et al., 2019; Fang et al., 2020; Shejwalkar and Houmansadr, 2021), we consider the goal of an attacker is to poison the training process of FL such that the final learnt global model makes incorrect predictions for a large number of clean inputs indiscriminately. In other words, the poisoned global model has a large testing error rate. Such a global model would eventually be abandoned, leading to denial-of-service attacks (Barreno et al., 2006).
@@
Attacker’s capability:  We assume that the attacker is capable of injecting fake clients into an FL system as malicious clients. We note that this is a practical threat model. Multiple platforms (sim, [n. d.]; Nox, [n. d.]; blu, [n. d.]) can be used to automatically inject fake devices/clients in a distributed system. For instance, when the FL system learns a global model on Android smartphones (i.e., clients), e.g., Google’s next-word prediction on Android devices (gbo, [n. d.]), an attacker can use a laptop to simulate fake smartphones via Android emulators (Cao and Gong, 2022).
@@
We note that our threat model is the same as previous fake client based model poisoning attacks (Cao and Gong, 2022), which do not require genuine clients’ information.
4. Motivation
@@
Figure 1. Illustration of (a) limited attack effectiveness and (b) inconsistent malicious model updates of existing attacks.
We first introduce our key observation about why existing model poisoning attacks achieve suboptimal attack effectiveness, which motivates the design of our PoisonedFL in the next section. We show experimental results on the MNIST dataset with the default parameter settings described in Section 6.1. We assume the server uses Trimmed Mean (Yin et al., 2018) as the defense. Moreover, we assume the ratio of fake clients to genuine clients is $20\%$. We consider MPAF (Cao and Gong, 2022), an attack not requiring genuine clients’ information, and Fang (Fang et al., 2020), a representative attack that requires genuine clients’ information. To give advantages to Fang, we assume the attacker knows the local models of all participating genuine clients in each round when crafting malicious model updates on fake clients.

### v2

# Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

@@ Abstract
1 Introduction
Federated Learning (FL) enables clients–such as smartphones, self-driving cars, and IoT devices–to collaboratively train a model without sharing their raw local training data with a central server [1, 2]. Typically, FL involves multiple training rounds, each of which consists of three key steps: 1) the server sends the current global model to the clients or a subset of them; 2) a client trains a local model based on the current global model and its local training data, and sends the model update (i.e., the difference between local model and global model) to the server; and 3) the server aggregates the clients’ model updates according to an aggregation rule and adds the aggregated model update to the global model. FL has wide applications across diverse domains, including healthcare [3, 4, 5] and finance [6, 7].
However, due to the distributed nature, FL is fundamentally vulnerable to model poisoning attacks [8, 9, 10, 11]. In particular, an attacker can control some malicious clients and send carefully crafted malicious model updates to the server to poison the global model. The malicious clients could be compromised genuine clients or injected fake clients. The poisoned global model misclassifies many inputs indiscriminately, i.e., it has a large testing error rate (also known as untargeted model poisoning attacks).
Existing model poisoning attacks suffer from two key limitations. First, they [8, 10, 11, 9] only leverage in-round consistency, i.e., they send the same or similar malicious model updates in a single training round to amplify the attack effect, but neglect multi-round consistency. For instance, Fang attack [10] crafts the malicious model updates in a given training round to maximally deviate the global model in that round, while MPAF [9] sets the malicious model updates in a training round as the magnified difference between the current global model and a random model. Such a strategy limits the attack effectiveness when defenses are deployed. This is because defenses weaken the attack effect in every training round by filtering and clipping the malicious model updates; and the weakened attack effects further self-cancel across training rounds due to the lack of multi-round consistency, as shown by our measurement results in Section 4. Second, most existing attacks [8, 10, 11] require knowledge of genuine clients’ model updates or local training data. They often assume compromising a significant number of genuine clients to obtain such knowledge, which is notably impractical. These attacks with suboptimal effectiveness and/or impractical assumptions give a false sense of FL security.
@@ 1 Introduction
We evaluate our PoisonedFL on five benchmark datasets. Our results show that, under the least knowledge scenario – that is, no knowledge about the defense nor genuine local training data/models – PoisonedFL can break eight state-of-the-art FL defenses [12, 13, 14, 15, 16, 17, 18]. For instance, PoisonedFL can reduce the global models to random guessing in many cases. Moreover, in most cases, PoisonedFL achieves higher testing error rates than existing attacks [8, 10, 11, 9] when a defense is deployed even if we assume the attacker has access to the local models of all genuine clients. In the remaining cases, both PoisonedFL and some existing attacks reduce the global models to random guessing. Furthermore, PoisonedFL remains effective across different FL settings including the fraction of malicious clients, the degree of non-IID of the genuine clients’ local training data, the number of epochs used to train a local model, and the fraction of clients selected to participate in each training round.
@@
2 Background and Related Work
2.1 Federated Learning (FL)
@@
Federated Learning (FL) [1, 2] is a machine learning paradigm that allows many clients (e.g., smartphones) to collaboratively train a model without sending their raw local training data to a server. Suppose we have $n$ clients and each client has a local training dataset $\mathcal{D}_{i}$, $i=1,2,\cdots,n$. FL aims to solve an optimization problem that is defined as $\min\sum_{i=1}^{n}\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$, where $\bm{w}\in\mathbb{R}^{d}$ is the model (called global model), $d$ is the number of parameters in the model $\bm{w}$, and $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ is the local loss function for client $i$. FL iteratively trains the global model $\bm{w}$. In the $t$-th training round, the server first distributes the current global model $\bm{w}^{t-1}$ to all clients or a subset of them. For simplicity, we assume all clients are selected, but we will also evaluate a subset of them are selected in each training round in our experiments. Each client $i$ then trains a local model $\bm{w}_{i}^{t}$ based on the current global model and its local training data. For instance, client $i$ initializes $\bm{w}_{i}^{t}$ as $\bm{w}^{t-1}$ and iteratively updates $\bm{w}_{i}^{t}$ via minimizing $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ using Stochastic Gradient Descent (SGD). After training the local model $\bm{w}_{i}^{t}$, client $i$ sends the model update $\bm{g}^{t}_{i}=\bm{w}^{t}_{i}-\bm{w}^{t-1}$ to the server. Finally, the server aggregates the received model updates and updates the global model as follows:
$\bm{w}^{t}=\bm{w}^{t-1}+\bm{g}^{t},$ (1)
@@
where $\bm{g}^{t}={{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the aggregated model update, ${AR}$ denotes the aggregation rule used by the server, and $\{\bm{g}^{t}_{i}\}_{i\in[1,n]}$ is the set of clients’ model updates. For instance, when ${AR}$ is FedAvg [2], the aggregated model update ${{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the average of the clients’ model updates.
2.2 Poisoning Attacks to FL
@@
In model poisoning attacks [19, 8, 10, 11], malicious clients corrupt the global model via sending carefully crafted malicious model updates to the server. In a targeted attack [19], a poisoned global model only misclassifies attacker-chosen inputs with certain properties (e.g., embedded with an attacker-chosen trigger). In contrast, in an untargeted attack [8, 10, 11], a poisoned global model indiscriminately misclassifies a large number of clean inputs, leading to a large testing error rate and eventually denial-of-service attacks. Therefore, we focus on untargeted attacks in this work.
@@
Requiring genuine clients’ information:  One category of model poisoning attacks [8, 10, 11] require that an attacker has access to genuine clients’ model updates or local training data. To obtain such knowledge, these attacks often assume an attacker can compromise a large fraction of genuine clients. For instance, in a little is enough (LIE) attack [8], the attacker computes the average of the genuine model updates of the compromised genuine clients and adds perturbation to them to craft the malicious model updates in each training round. Concurrent to LIE, Fang et al. [10] proposed the first optimization-based attack framework that can be tailored to any aggregation rule. Their key idea is that the attacker crafts the malicious model updates such that the aggregated model update after attack deviates substantially from the before-attack one in each training round. Building upon this attack framework [10], the authors in [11] assume the malicious model update to be the sum of the before-attack aggregated model update and a fixed perturbation vector scaled by a factor, where the scaling factor is optimized to maximize the deviation between the after-attack aggregated model update and the before-attack one.
@@
Not requiring genuine clients’ information:  The other category of model poisoning attacks [9] do not rely on any genuine clients’ information. Therefore, they only assume that an attacker can inject fake clients into an FL system, which is more realistic than compromising genuine clients. Since genuine clients’ information is not available, these attacks craft malicious model updates based on heuristics. For instance, in MPAF [9], the attacker crafts malicious model updates such that the global model progressively moves towards a random model in each training round.
@@
Limitations of existing attacks:  Existing attacks suffer from the following limitations. First, the first category of attacks require an attacker to compromise a large fraction of genuine clients, which is not affordable in production FL [20]. Second, both categories of attacks craft malicious model updates independently in different training rounds, leading to self-cancellation of attack effect across training rounds. As a result, they achieve suboptimal attack effectiveness when defenses are deployed.
2.3 Defenses against Poisoning Attacks
@@
Byzantine-robust aggregation rules:  Instead of using the average of the clients’ model updates to update the global model, Byzantine-robust aggregation rules [12, 13, 21, 15, 16, 17, 22, 23] typically employ statistical analysis to filter out or clip outlier model updates before aggregating them. For instance, Multi-Krum [12] iteratively selects the model update that is the closest to its nearest neighbor model updates, and aggregates them. Dimension-wise Median (or Trimmed Mean) [17] uses the median (or trimmed-mean) of the clients’ local models for each dimension. Norm Bound [16] aims to minimize the impact of potential malicious model updates by clipping the magnitudes of all model updates. FLTrust [13] assumes the server uses a clean and small dataset to bootstrap trust. Specifically, in FLTrust [13], the server trains a server model and assigns trust scores to clients by measuring the similarity between the server model update and clients’ model updates. Under some assumptions about the learning problem and clients’ local training data distribution, these Byzantine-robust aggregation rules often can theoretically guarantee that the learnt global model does not change much when the number of malicious clients is bounded. However, they are still vulnerable to our attack because these assumptions often do not hold and a bounded change in the parameters of the global model does not guarantee a small change of error rate.
@@
Provably robust defenses:  Provably robust defenses [24, 14, 25] can provide stronger theoretical guarantees than Byzantine-robust aggregation rules. In particular, given a number of malicious clients, a provably robust defense can guarantee a lower bound of testing accuracy no matter what poisoning attacks the malicious clients use. For instance, FLCert [14] trains multiple global models, each of which is trained using a subset of clients. Given a testing input, FLCert predicts its label by taking a majority vote among the multiple global models. A major limitation of these provably robust defenses is that they can only tolerate a small number of malicious clients, i.e., the lower bound of testing accuracy reduces to 0 when the number of malicious clients is moderate. As a result, they are still vulnerable to our attack since an attacker can easily inject a large number of fake clients.
@@
Malicious clients detection:  These defenses [26, 18] aim to detect malicious clients during or after the training process. After detecting malicious clients, the server can further efficiently recover/re-train an accurate global model using the remaining clients [27]. For instance, FLDetector [18] aims to detect malicious clients by utilizing the inconsistency of a client’s model updates across multiple training rounds. Their key idea is that the model updates of a benign client in multiple training rounds are all calculated based on the client’s local training data, while the model updates of a malicious client are crafted without following the learning procedure. As a result, the model updates of a benign client satisfy certain consistency constraints across training rounds, while those of a malicious client do not, making it possible to distinguish between benign and malicious clients. In our attack, we enforce the malicious model updates of a malicious client to be consistent across training rounds, making FLDetector ineffective.
3 Threat Model
@@
Attacker’s goal:  Following previous works  [8, 10, 11], we consider the goal of an attacker is to poison the training process of FL such that the final learnt global model makes incorrect predictions for a large number of clean inputs indiscriminately. In other words, the poisoned global model has a large testing error rate. Such a global model would eventually be abandoned, leading to denial-of-service attacks [28].
@@
Attacker’s capability:  We assume that the attacker is capable of injecting fake clients into an FL system as malicious clients. We note that this is a practical threat model. Multiple platforms [29, 30, 31] can be used to automatically inject fake devices/clients in a distributed system. For instance, when the FL system learns a global model on Android smartphones (i.e., clients), e.g., Google’s next-word prediction on Android devices [32], an attacker can use a laptop to simulate fake smartphones via Android emulators [9].
@@
We note that our threat model is the same as previous fake client based model poisoning attacks [9], which do not require genuine clients’ information.
4 Motivation
@@
Figure 1: Illustration of (a) limited attack effectiveness and (b) inconsistent malicious model updates of existing attacks.
We first introduce our key observation about why existing model poisoning attacks achieve suboptimal attack effectiveness, which motivates the design of our PoisonedFL in the next section. We show experimental results on the MNIST dataset with the default parameter settings described in Section 6.1. We assume the server uses Trimmed Mean [17] as the defense. Moreover, we assume the ratio of fake clients to genuine clients is $20\%$. We consider MPAF [9], an attack not requiring genuine clients’ information, and Fang [10], a representative attack that requires genuine clients’ information. To give advantages to Fang, we assume the attacker knows the local models of all participating genuine clients in each round when crafting malicious model updates on fake clients.